\begin{thebibliography}{00}

\bibitem{ADN64} S. Agmon, A. Douglis, and L. Nirenberg, Estimates near the boundary for solutions of elliptic partial differential equations satisfying general boundary conditions II, Commun. Pure Appl. Math., 17 (1964), 35--92.

\bibitem{cheungsuen16} K.L. Cheung and A. Suen, Existence and uniqueness of small energy weak solution to multi-dimensional compressible Navier-Stokes equations with large external potential force, Journal of Mathematical Physics, Vol. 57 (2016), no. 8, 081513, 1--19.

\bibitem{danchin00} R. Danchin, Global existence in critical spaces for compressible Navier-Stokes equations, Inventiones Mathematicae, 141(3) (2000), 579--614.

\bibitem{danchin01} R. Danchin, Global existence in critical spaces for flows of compressible viscous and heat conductive gases, Arch. Ration. Mech. Anal., 160 (2001), 1--39.

\bibitem{evans10} L. C. Evans, Partial Differential Equations, Grad. Texts in Math. 19, American Mathematical Society, Providence, RI, 2010.

\bibitem{feireisl04} E. Feireisl, Dynamics of Viscous Compressible Fluids, Oxford Lecture Series in Mathematics and its Applications, 26. Oxford University Press, Oxford, 2004. 

\bibitem{hoff95} D. Hoff, Global solutions of the Navier-Stokes equations for multidimensional, compressible flow with discontinuous initial data, J. Diff. Eqns. 120, no. 1 (1995), 215--254.

\bibitem{hoff05} D. Hoff, Compressible Flow in a Half-Space with Navier Boundary
Conditions, J. Math. Fluid Mech.  7 (2005), 315--338.

\bibitem{hoff06} D. Hoff, Uniqueness of weak solutions of the Navier-Stokes equations of multidimensional compressible flow, SIAM J. Math. Anal. 37 (6) (2006), 1742--1760.

\bibitem{hoff11} D. Hoff, Local solutions of a compressible flow problem with Navier boundary conditions in general three-dimensional domains. Siam J. Math. Anal. 44 (2011), 1872--1886.

\bibitem{hoffperepelitsa12} D. Hoff and M. Perepelitsa, Boundary tangency for density interfaces in compressible viscous flows, J. Differential Equations, 253 (2012), 12, 3543--3567.

\bibitem{hoffserre91} D. Hoff, D. Serre, The failure of continuous dependence on initial data for the Navier-Stokes equations of compressible flow, SIAM J. Appl. Math. 51 (1991) 887--898.

\bibitem{huwang08} Hu, X. and Wang, D.,
{\it Global solutions to the three-dimensional full compressible magnetohydrodynamic flows},
Comm. Math. Phys. 283, no. 1 (2008), 255--284.

\bibitem{kelliher06} J. Kelliher, Navier–Stokes equations with Navier boundary conditions for a bounded domain in the plane. SIAM J. Math. Anal. 38(1), 210--232 (2006)

\bibitem{kim09} N. Kim, Large friction limit and the inviscid limit of 2D Navier-Stokes equations under Navier friction condition, SIAM J. Math. Anal., 41 (2009), 1653--1663.

\bibitem{lions98} P.L. Lions, Mathematical Topics in Fluid Mechanics, vol. 2,
Oxford Lecture Series in Mathematics, 10 (1998).

\bibitem{mn79} A. Matsumura and T. Nishida, The initial value problem for the equation of
motion of compressible viscous and heat-conductive fluids, Proc. Japan. Acad.,
55 (1979), Ser. A, 337--342.

\bibitem{mn80} A. Matsumura and T. Nishida, The initial value
problem for the equations of motion of viscous and heat-conductive
gases, J. Kyoto Univ. 20 (1980), 67--104.

\bibitem{mn82} A. Matsumura and T. Nishida, Initial boundary value problems for the equations of motion of general fluids, in Computing Methods in Science and Engineering V, R. Glowinski and J.L. Lions, eds., North-Holland, Amsterdam, 1982.

\bibitem{Navier23} M. Navier, Sur les Lois du Mouvement des Fluides, M\'{e}moires de L'Acad\'{e}mie Royale des Sciences, Paris, Tome VI, 1823.

\bibitem{perepelitsa14} M. Perepelitsa, Weak solutions of the Navier-Stokes equations for compressible flows with no-slip boundary conditions, Arch. Rational Mech. Anal. 212 3 (2014), 709--726.

\bibitem{suenhoff12} A. Suen and D. Hoff, Global low-energy weak solutions of the equations of 3D compressible magnetohydrodynamics, Arch. Rational Mechanics Ana., 205 (2012), 27--58.

\bibitem{suen13} A. Suen, Global Solutions of the Navier-Stokes Equations for Isentropic Flow with Large External Potential Force, Z. Angew. Math. Phys. (ZAMP), Vol. 64 (2013), no. 3, 767--784.

\bibitem{suen14} A. Suen, Existence of global weak solution to Navier-Stokes equations with large external potential force and general pressure, Math. Methods Appl. Sci., Vol. 37, no. 17 (2014), 2716--2727.

\bibitem{suen20a} A. Suen, Existence and uniqueness of low-energy weak solutions to the compressible 3D magnetohydrodynamics equations, J. Diff. Eqns, Vol. 268 (6) (2020), 2622--2671.

\bibitem{suen20b} A. Suen, Existence and a blow-up criterion of solution to the 3D compressible Navier-Stokes-Poisson equations with finite energy, Discrete and Continuous Dynamical Systems Series A (DCDS-A), Vol. 40 (2020), no. 3, 1775--1798.

\bibitem{tani77} A. Tani, On the first initial-boundary value problem of compressible viscous fluid motion, Publ. Res. Inst. Math. Sci. 13 (1977), 193--253.

\bibitem{wehrheim04} K. Wehrheim, Uhlenbeck Compactness, EMS Ser. Lectures in Math., Z\"{u}rich, 2004.

\bibitem{ziemer89} W. Ziemer, Weakly differentiable functions, Springer-Verlag, 1989.
\end{thebibliography}

\documentclass[reqno]{amsart}
 \theoremstyle{definition}
 
 \theoremstyle{remark}

 \numberwithin{equation}{section}

\usepackage{graphicx}
\usepackage{amssymb, amsmath, amsthm}
\usepackage{amsfonts}
\usepackage{amssymb}
\usepackage{float}
\usepackage{mathrsfs}
\usepackage{hyperref}
\hypersetup{
    bookmarks=true,         
    pdffitwindow=false,     
    pdfstartview={FitH},    
    colorlinks=false,       
}
\usepackage{enumitem} 
\newtheorem{theorem}{Theorem}

\newtheorem{definition}[theorem]{Definition}

\newtheorem{lemma}[theorem]{Lemma}

\newtheorem{remark}[theorem]{Remark}

\newcommand{\eqnum}{\refstepcounter{equation}\textup{\tagform@{\theequation}}}
\newcommand{\dis}{\displaystyle}

\newcommand{\divv}{\text{\rm div}}

\newcommand{\No}{{\bf n}}
\newcommand{\Ta}{{\bf{t}}}

\newcommand{\cL}{\mathcal L}

\newcommand{\R}{\mathbb R}

\newcommand{\intoxb}{\int_{\partial\Omega}}
\newcommand{\intox}{\int_{\Omega}}
\newcommand{\intoxt}{\int_0^t\int_{\Omega}}
\newcommand{\intoxtb}{\int_0^t\int_{\partial\Omega}}
\newcommand{\trho}{\tilde\rho}
\newcommand{\tu}{\tilde{\bf u}}
\newcommand{\vu}{{\bf u}}
\newcommand{\vw}{{\bf w}}
\newcommand{\seta}{\tilde\sum_1^{M,T}}
\newcommand{\setb}{\tilde\sum_2^{N,T}}

\renewcommand{\theequation}{\thesection.\arabic{equation}}
\numberwithin{equation}{section}
\numberwithin{theorem}{section}

\numberwithin{figure}{section}

\begin{document}

%
%
%
%
%
%
%
%
%








\title[Large friction limit in general domains]
 {Large friction limit of the compressible Navier-Stokes equations with Navier Boundary conditions in general three-dimensional domains}
 
\author{Anthony Suen} 

\address{Department of Mathematics and Information Technology\\
The Education University of Hong Kong}

\email{acksuen@eduhk.hk}

\date{July 1, 2020}

\keywords{Navier-Stokes equations; compressible flow; Navier boundary conditions; large friction limit}

\subjclass[2000]{35Q30} 

\begin{abstract}
In this paper, we study the Navier-Stokes equations of compressible, barotropic flow posed in a bounded set in $\R^3$ with different boundary conditions. Specifically, we prove that the local-in-time smooth solution of the Navier-Stokes equations with Navier boundary condition converges to the smooth solution of the Navier-Stokes equations with no-slip boundary condition as the Navier friction coefficient tends to infinity. 
\end{abstract}

\maketitle
\section{Introduction}\label{Introduction section}

We prove that the local-in-time smooth solution of the Navier-Stokes equations with Navier boundary condition converges to the smooth solution of the Navier-Stokes equations with no-slip boundary condition as the Navier friction coefficient tends to infinity. The present work is intended as the first step in extending the convergence results of large friction limit from the incompressible flows to compressible flows.

The Navier-Stokes equations of a compressible, barotropic flow in a bounded domain $\Omega\subset \R^3$ give the conservation of mass and the balance of momentum:
\begin{align}
\label{Navier Stokes} \left\{ \begin{array}{l}
\rho_t + \divv (\rho \vu) =0, \\
(\rho u^j)_t + \divv (\rho u^j \vu) + (P(\rho))_{x_j} = \mu\,\Delta u^j +
\lambda \, (\divv \,\vu)_{x_j}.
\end{array}\right.
\end{align}
The above system \eqref{Navier Stokes} is solved together with initial conditions
\begin{align}\label{initial condition}
(\rho(\cdot,0),\vu(\cdot,0))=(\rho_0,\vu_0)
\end{align}
and is equipped with {\it either one} of the following boundary conditions, namely
\begin{align}\label{Navier condition}
\mbox{$\No\cdot \vu(x)=0$ and $K\mu(\nabla \vu\No)\cdot \Ta(x)=- \vu\cdot \Ta(x)$, \qquad for $x\in\partial\Omega$}
\end{align}
or
\begin{align}\label{no slip condition}
\vu(x)=0,\qquad\mbox{for $x\in\partial\Omega$}.
\end{align}
The meanings for the functions and symbols are given as follows:
\begin{itemize}
\item $\rho$ and $\vu=(u^1,u^2,u^3)$ are functions of $x\in\Omega$ and $t\ge0$ which represent density and velocity respectively;
\item $P = P(\rho)$ is a given function in $\rho$ which stands for the pressure;
\item $\mu$, $\lambda > 0$ are viscosity constants;
\item $(\cdot)_{x_j}$ and $(\cdot)_t$ stand for the spatial derivative $\frac{\partial}{\partial x_j}$ and time derivative $\frac{\partial}{\partial t}$ respectively;
\item $\divv(\cdot)$ and $\Delta(\cdot)$ are the usual spatial divergence and Laplace operators;
\item $\No(x)$ and $\Ta(x)$ are the unit outward normal and tangent vectors respectively on $\partial \Omega$;
\item $K>0$ is a constant and $\alpha:=\frac{1}{K}$ is known as the {\it Navier friction coefficient}.
\end{itemize}
To facilitate later discussions, we name the system \eqref{Navier Stokes}-\eqref{initial condition} with boundary condition \eqref{Navier condition} the {\it Navier-Stokes equations with Navier boundary condition} (NSENC), while we name the system \eqref{Navier Stokes}-\eqref{initial condition} with boundary condition \eqref{no slip condition} the {\it Navier-Stokes equations with Dirichlet boundary condition} (NSEDC).

The so-called Navier boundary condition \eqref{Navier condition} was first proposed by Navier in \cite{Navier23} which states that the velocity on $\partial\Omega$ is proportional to the tangential component of the stress, while the Dirichlet boundary condition \eqref{no slip condition} (or more precisely the {\it no-slip} boundary condition) assumes the fluid will have zero velocity relative to the boundary. Regarding the two different boundary conditions given by \eqref{Navier condition} and \eqref{no slip condition}, a very interesting but natural question arises:
\begin{itemize}
\item[$(\mathcal{Q})$:] {\it As the constant $K$ vanishes (or equivalently the Navier friction coefficient $\alpha$ tends to infinity), will the solutions to the Navier-Stokes equations with Navier boundary conditions converge to a solution to the Navier-Stokes equations with the usual no-slip boundary conditions?}
\end{itemize}
The present work is devoted to answer the question $(\mathcal{Q})$ by justifying the limit for smooth local-in-time solutions of \eqref{Navier Stokes}-\eqref{initial condition} with boundary condition \eqref{Navier condition} as $K$ tends to zero.

We first recall some known existence results regarding the system \eqref{Navier Stokes}-\eqref{initial condition}. Generally speaking, by imposing different conditions on the initial data, three types of solutions can be shown to exist and they enjoy different properties:
\begin{itemize}
\item[(i)] When the initial data is taken to be close to a constant in $H^3(\Omega)$, Matsumura and Nishida \cite{mn79, mn80, mn82} obtained global-in-time ``small-smooth type'' solutions, which were later extended by Danchin \cite{danchin00, danchin01} to solutions in certain scale-invariant homogeneous Besov spaces. Small-smooth type solutions are constructed via iterative procedure based on asymptotic decay rates for the corresponding linearised system, and they do not exhibit the generic singularities of the system. 
\item[(ii)] When the initial data is having arbitrarily large energy and with nonnegative density, Lions \cite{lions98} and Feireisl \cite{feireisl04} proved the existence of ``large-weak type'' solutions (also refer to \cite{huwang08} for improvements in such direction). Large-energy weak solutions by their very nature possess very little regularity, which may even include some non-physical solutions (see \cite{hoffserre91} for related discussions). 
\item[(iii)] Different from the two types of solutions as mentioned in (i) and (ii), a third type of ``intermediate-weak type'' solutions were studied by Hoff \cite{hoff95, hoff05}, Perepelitsa \cite{perepelitsa14}, Suen \cite{suen13, suen14, suen20b} and Suen-Hoff \cite{suenhoff12} for which initial data are small in fairly weak norms and initial densities are nonnegative and essentially bounded. Such intermediate-weak type solutions have rich physical and mathematical meanings compared to other solution classes: on the one hand, these solutions may exhibit discontinuities in density and velocity gradient across hypersurfaces in $\R^3$, which is not observable from small-smooth solutions; on the other hand, the solutions would still have enough regularity for the development of a uniqueness and continuous dependence theory \cite{cheungsuen16, hoff06, suen20a} which seems difficult from the very weak framework for large-weak type solutions. Moreover, such solutions demonstrate fine-structure property near the boundary under no-slip boundary condition \eqref{no slip condition} (refer to \cite{hoffperepelitsa12} for the instantaneous tangency for density interfaces), which behave drastically different from the case when the Navier boundary condition \eqref{Navier condition} is imposed (see \cite{hoff05}).
\end{itemize}
Our goal is to first address the convergence of smooth local-in-time solutions (which can be viewed as the category (i) solutions mentioned above but without the smallness assumption on the initial data), which is the central topic for this paper. The convergences of weak solutions from category (ii) and (iii) will be considered elsewhere.

The convergence issue raised by question $(\mathcal{Q})$ was studied by Kelliher \cite{kelliher06} and Kim \cite{kim09} for incompressible Navier-Stokes equations which proved the convergence of smooth solutions as $K$ vanishes. In contrast to the incompressible flows, the compressible system \eqref{Navier Stokes} contains the unknown density function $\rho$, which contributes to the following difficulties:
\begin{itemize}
\item one has to obtain {\it appropriate} bounds on $\rho$ in order to gain control over the velocity $\vu$, which is different from the incompressible case for which one can automatically have some control over $\vu$ on the boundary (see \cite[Lemma~9.1]{kelliher06});
\item one has to ensure some {\it $K$-independent} bounds on both $\rho$ and $\vu$ before taking the limit $K\to0$, which is highly non-trivial compared with the incompressible cases.
\end{itemize} 
In this present work, we try to extend the results from \cite{kelliher06} and \cite{kim09} to the case for compressible Navier-Stokes equations and obtain convergence of {\it both} densities and velocities. The main novelties are:
\begin{itemize}
\item We obtain bounds on local-in-time solutions to \eqref{Navier Stokes}-\eqref{initial condition} with boundary condition \eqref{Navier condition} which are {\it independent of} $K$;
\item We prove the strong convergence of both densities and velocities as $K$ vanishes, which extends and strengthen the results for incompressible Navier-Stokes equations.
\end{itemize} 

We now give a precise formulation of our results. The parameters $\Omega$, $\mu$, $\lambda$, $P$ and $K$ will be assumed to satisfy the following:
\begin{align}\label{condition on the domain bdry}
\mbox{$\Omega$ is a bounded open set in $\R^d$ with a $C^4$ boundary;}
\end{align}
\begin{align}\label{conditions on viscosity}
\mbox{$\mu>0$ and $\mu+3\lambda>0$;}
\end{align}
\begin{align}\label{condition on pressure}
\mbox{$P\in C([0, \infty))\cap C^2((0,\infty));$}
\end{align}
\begin{align}\label{condition on K}
\mbox{$K$ is independent of $x$ and $t$ with $K>0$.}
\end{align}
Concerning the initial data $(\rho_0,\vu_0)$, it will be assumed that there is a positive constant $M_0$ such that
\begin{align}\label{boundedness condition on initial data}
\|\rho_0\|_{L^\infty},\|\rho_0^{-1}\|_{L^\infty},\|\rho_0\|_{H^2},\|\vu_0\|_{H^3}\le M_0,
\end{align}
and $\vu_0$ further satisfies
\begin{align}\label{compatibility condition on u0 no-slip}
\vu_0(x)=0,\qquad\mbox{for $x\in\partial\Omega$}.
\end{align}

We have the following local-in-time existence theorem for smooth solutions to \eqref{Navier Stokes}. It gives the necessary bounds on the smooth solutions and will be useful for proving the convergence results later. The proof will be given in Section~\ref{existence section}. 

\begin{theorem}\label{local-in-time existence theorem}
Assume that the hypotheses \eqref{condition on the domain bdry}-\eqref{condition on K} hold and let positive
numbers $M_0$ and $M_0'>M_0$ be given. Then there is a positive time $T>0$ and a constant $N$, both depending on $\Omega$, $\mu$, $\lambda$, $P$ $M_0$ and $M_0'$ but independent of $K$ such
that if initial data $(\rho_0,\vu_0)$ is given satisfying \eqref{boundedness condition on initial data}-\eqref{compatibility condition on u0 no-slip}, then there are solutions $(\rho, \vu)$ and $(\trho,\tu)$ to the initial-boundary value problems (NSENC) and (NSEDC) respectively which are defined on $\Omega \times [0, T]$ and they satisfy
\begin{align}\label{boundedness on rho local-in-time solution}
&\sup_{0\le t\le T} (\|\rho(\cdot,t)\|_{L^\infty},\|\rho^{-1}(\cdot,t)\|_{L^\infty},\|\rho(\cdot,t)\|_{H^2})\notag\\
&\qquad+\sup_{0\le t\le T} (\|\trho(\cdot,t)\|_{L^\infty},\|\trho^{-1}(\cdot,t)\|_{L^\infty},\|\trho(\cdot,t)\|_{H^2})\le M_0'
\end{align}
and
\begin{align}\label{boundedness on u local-in-time solution}
&\sup_{0\le t\le T} (\|\vu(\cdot,t)\|_{H^3}^2,\|\vu_t(\cdot,t)\|_{H^1}^2)+\int_0^T\|\vu_t(\cdot,t)\|_{H^2}^2dt\notag\\
&\qquad+\sup_{0\le t\le T} (\|\tu(\cdot,t)\|_{H^3}^2,\|\tu_t(\cdot,t)\|_{H^1}^2)+\int_0^T\|\tu_t(\cdot,t)\|_{H^2}^2dt\le N.
\end{align}
\end{theorem}

\begin{remark}
Here are some remarks regarding Theorem~\ref{local-in-time existence theorem}
\begin{itemize}
\item In view of Theorem~\ref{local-in-time existence theorem}, for the case of (NSENC), one can replace the compatibility condition \eqref{compatibility condition on u0 no-slip} on $\vu_0$ with a weaker assumption, namely
\begin{align*}
\mbox{$\No\cdot \vu_0(x)=0$ and $K\mu(\nabla \vu_0\No)\cdot \Ta(x)=- \vu_0\cdot \Ta(x)$, \qquad for $x\in\partial\Omega$.}
\end{align*}
\item The key idea of proving Theorem~\ref{local-in-time existence theorem} follows closely to the one given in \cite{hoff11}, the main difference here is to show the $K$-independence of the constants $M'_0$ and $N$.
\item It is important to have $M'_0$ and $N$ both being independent of $K$, which allow us to apply the bounds \eqref{boundedness on rho local-in-time solution}-\eqref{boundedness on u local-in-time solution} for proving the convergence of smooth solutions as $K$ vanishes.
\end{itemize}
\end{remark}

Once we obtain the local-in-time existence of smooth solutions to to the systems (NSENC) and (NSEDC), we proceed to study the convergence as $K$ vanishes. The following theorem is the main result of this paper.
\begin{theorem}\label{main thm}
Assume that the hypotheses \eqref{condition on the domain bdry}-\eqref{condition on K} hold and let $(\rho_0,\vu_0)$ be given functions which satisfy \eqref{boundedness condition on initial data}-\eqref{compatibility condition on u0 no-slip}. Suppose $(\rho,\vu)$ and $(\trho,\tu)$ are smooth local-in-time solutions to the systems (NSENC) and (NSEDC) respectively defined on $\Omega\times[0,T]$ with the same initial data $(\rho_0,\vu_0)$, as described by Theorem~\ref{local-in-time existence theorem}. Then there exists $T^*\in(0,T]$ such that for any $s_1\in[0,2]$ and $s_2\in[0,1]$,
\begin{align}\label{convergence of u in Hs}
\vu\to\tu\mbox{ in $L^\infty([0,T^*];H^{s_1}(\Omega))$ as $K\to0$,}
\end{align}
\begin{align}\label{convergence of u on the boundary}
\vu\to\tu\mbox{ in $L^2([0,T^*];L^2(\partial\Omega))$ as $K\to0$,}
\end{align}
and
\begin{align}\label{convergence of rho in Hs}
\rho\to\trho\mbox{ in $L^\infty([0,T^*];H^{s_2}(\Omega))$ as $K\to0$.}
\end{align}
\end{theorem}

\begin{remark}
As a consequence of Theorem~\ref{main thm}, if we define the energy functional $E(\rho,\vu,t)$ by
\begin{align*}
E(\rho,\vu,t)=\sup_{0\le s\le t}\left[\int_{\Omega}(|\vu|^2+|\rho|^2)(x,s)dx\right]+\int_0^t\int_{\Omega}|\nabla \vu|^2(x,s)dxds,
\end{align*}
then for all $t\in[0,T^*]$, we have $E(\rho,\vu,t)\to E(\trho,\tu,t)$ as $K\to0$. 
\end{remark}

The rest of the paper is organised as follows. In Section~\ref{prelim section}, we give some preliminary facts and  definitions used in this paper and provide an estimate for the Lam\'{e} operator defined in \eqref{def of lame}. In Section~\ref{existence section}, we obtain bounds on $\rho$ and $\vu$ which are independent of $K$ and prove Theorem~\ref{local-in-time existence theorem}. Finally in Section~\ref{convergence section}, we prove the convergence of smooth solutions to \eqref{Navier Stokes} as $K$ vanishes by making use of the bounds obtained in Theorem~\ref{local-in-time existence theorem}, thereby proving Theorem~\ref{main thm}. 

\section{Preliminaries and notations}\label{prelim section}

We introduce the following notations and conventions:
\begin{itemize}
\item For $s\ge0$ and $p\ge1$, $W^{s,p}(\Omega)$ is the usual Sobolev space with norm $\|\cdot\|_{W^{s,p}(\Omega)}$ given by
\begin{align*}
\|\cdot\|_{W^{s,p}(\Omega)}(f):=\sum_{|\beta|\le s} \Big(\intox|D^{\beta}_x f(x)|^pdx\Big)^\frac{1}{p}.
\end{align*}
We write $H^s(\Omega):=W^{s,2}(\Omega)$. For simplicity, we also write $H^s=H^s(\Omega)$, $\|\cdot\|_{L^p}=\|\cdot\|_{L^p(\Omega)}$, $\|\cdot\|_{W^{s,p}}=\|\cdot\|_{W^{s,p}(\Omega)}$, etc. unless otherwise specified.
\item We adopt the usual notation for H\"older seminorms, namely for $v:\R^3\to \R^3$ and $\beta \in (0,1]$, 
\hfill
$$\langle v\rangle^\beta = \sup_{{x_1,x_2\in 
\R^3}\atop{x_1\not=x_2}}
{{|v(x_2) -v(x_1)|}\over{|x_2-x_1|^\beta}}\,;$$
and for $v:Q\subseteq\R^3 \times[0,\infty)\to \R^3$ and $\beta_1,\beta_2 \in (0,1]$,
\hfill
$$\langle v\rangle^{\beta_1,\beta_2}_{Q} = \sup_{{(x_1,t_1),(x_2,t_2)\in 
Q}\atop{(x_1,t_1)\not=(x_2,t_2)}}
{{|v(x_2,t_2) - v(x_1,t_1)|}\over{|x_2-x_1|^{\beta_1} + |t_2-t_1|^{\beta_2}}}\,.$$
\item Regarding the constants used in this work, $C$ shall denote a positive and sufficiently large constant, whose value may change from line to line.
\end{itemize}

We recall the following standard facts (see \cite{evans10} and \cite{ziemer89} for example) which will be useful for later analysis:
\begin{itemize}
\item There is a constant $C = C(\Omega)$ such that for all $\varphi\in H^2$,
\begin{align}\label{bound on L infty norm}
\|\varphi\|_{L^\infty}\le C\|\varphi\|_{H^2}.
\end{align}
\item There is a constant $C=C(\Omega)$ such that for $p\in[2,6]$ and $\varphi\in H^1$,
\begin{align}\label{embedding from Lp to H1}
\|\varphi\|_{L^p}\le C\Big(\|\varphi\|_{L^2}+\|\varphi\|_{L^2}^\frac{6-p}{2p}\|\nabla\varphi\|_{L^2}^\frac{3p-6}{2p}\Big).
\end{align}
\item Assume that $\partial\Omega$ is $C^1$, then there exists a constant $C = C(\Omega)$ such that for all $\varphi\in W^{1,1}$,
\begin{align}\label{bound on bdry integral}
\int_{\partial\Omega}|\varphi(x)|^2dS_x\le C\intox\{|\varphi(x)|^2+|\varphi(x)||\nabla\varphi(x)|\}dx.
\end{align}
\end{itemize}

For $\mu>0$ and $\mu+3\lambda>0$, we let $\cL$ denote the Lam\'{e} operator given by
\begin{align}\label{def of lame}
(\cL \vu)^j=\mu\Delta u^j + \lambda \divv(\vu_{x_j}).
\end{align}
With respect to the Lam\'{e} operator $\cL$, we define the system
\begin{align}
\label{Lame system} \left\{ \begin{array}{l}
\cL \vu = -g \mbox{\,\,\,in $\Omega$,}\\
\mbox{$\No\cdot \vu=0$ and $K\mu(\nabla \vu\No)\cdot \Ta=- \vu\cdot \Ta$ on $\partial\Omega$,}
\end{array}\right.
\end{align}
where $\vu:\bar{\Omega}\to\R^3$ is the unknown function and $g:\Omega\to\R^3$ is given. 

The following lemma gives an estimate on $\vu$ in terms of $g$ which is crucial for obtaining bounds on solutions to the system (NSENC) later.
 
\begin{lemma}\label{estimate on lame operator lemma}
Assume that $\mu>0$ and $\mu+3\lambda>0$ and let $m=0$ or 1. Assume in addition that $\Omega$ is a bounded open set in $\R^3$ with a $C^{m+3}$ boundary. Then there is a constant $C = C(\Omega,\mu,\lambda)$ independent of $K$ such that if $\vu$ is a solution of \eqref{Lame system} with $g \in H^m(\Omega)$, then $\vu \in H^{m+2}(\Omega)$ and
\begin{align}\label{elliptic estimate on lame operator}
\|\vu\|_{H^{m+2}}\le C(\|\vu\|_{L^2}+\|g\|_{H^m})
\end{align}
\end{lemma}

\begin{proof}
The proof is almost identical to the one given in \cite[Lemma~2.2]{hoff11} (also refer to \cite{ADN64} and \cite{wehrheim04} for more details). To see why the constant $C$ is independent of $K$, we give the estimates on $\|\nabla\vu\|_{L^2}$ as an example. We multiply \eqref{Lame system}$_1$ by $\vu$ and integrate to get
\begin{align}\label{H1 estimate on u lame system}
\intox\cL \vu\cdot \vu dx=-\intox g\cdot\vu dx.
\end{align}
Upon integrating by parts and applying the boundary condition \eqref{Lame system}$_2$, the integral on the left of \eqref{H1 estimate on u lame system} can be rewritten as follows
\begin{align*}
\intox\cL \vu\cdot \vu dx=-\alpha\intoxb|\vu|^2dS_x-\intox[\mu|\nabla\vu|^2+\lambda(\divv(\vu))^2]dx,
\end{align*}
where $\alpha=\frac{1}{K}$. Hence \eqref{H1 estimate on u lame system} becomes
\begin{align*}
\intox[\mu|\nabla\vu|^2+\lambda(\divv(\vu))^2]dx&=\intox g\cdot\vu dx-\alpha\intoxb|\vu|^2dS_x\\
&\le\intox g\cdot\vu dx,
\end{align*}
where the last inequity follows since $\alpha>0$. Therefore, we can see that $\|\nabla\vu\|_{L^2}$ can be bounded in terms of $\|g\|_{L^2}$ and is independent of $K$, which gives the required interior regularity estimates for $\vu$. The case when $\vu$ is supported in the intersection of $\Omega$ with a small neighborhood of a point on the boundary $\partial\Omega$ follows by the same argument given in \cite{hoff11} (which is somewhat simpler in our case here since $K$ is now just a positive constant) and we omit the details for the sake of brevity.
\end{proof}

\begin{remark}
We observe that $K$ (or equivalently $\alpha$) has the {\it correct} sign, which allows us to discard those corresponding boundary terms appeared in the above analysis. This is also important for proving the $K$-independence for the constant $C$ as in \eqref{elliptic estimate on lame operator}.
\end{remark}

\section{Existence of local-in-time smooth solutions: Proof of Theorem~\ref{local-in-time existence theorem}}\label{existence section}

In this section, we give the proof of Theorem~\ref{local-in-time existence theorem}. We only focus on the system (NSENC) since the local-in-time existence of (NSEDC) can be proved in the same way as in \cite{tani77}. Throughout this section $M_0$ will be fixed as in the statement of Theorem~\ref{local-in-time existence theorem}

We first give the following definitions of function spaces which will be useful for later analysis. More details can be found in \cite{hoff11}. 

\begin{definition}
For $M\ge M_0$ and $T>0$, $\seta$ is the set of maps $\rho:[0,T]\to H^2(\Omega)$ such that $\rho(\cdot,0)=\rho_0$, $\rho\in C([0,T];H^1(\Omega))$, $\rho_t\in C([0,T];L^2(\Omega))$, and 
\begin{align}\label{bound on rho def}
\sup_{0\le t\le T}(\|\rho(\cdot,t)\|_{L^\infty},\|\rho^{-1}(\cdot,t)\|_{L^\infty},\|\rho(\cdot,t)\|_{H^2})\le M.
\end{align}
And for $N>0$, $\setb$ is the set of maps $\vu:[0,T]\to H^3(\Omega)$ such that $\No\cdot \vu(x)=0$, $\vu(\cdot,0)=\vu_0$, $\vu\in C([0,T];H^2(\Omega))$, $\vu_t\in C([0,T];L^2(\Omega))$ and
\begin{align}\label{bound on u def}
\sup_{0\le t\le T}(\|\vu(\cdot,t)\|^2_{H^3},\|\vu_t(\cdot,t)\|^2_{H^1})+\int_0^T\|\vu_t(\cdot,t)\|^2_{H^2}\le N.
\end{align}
\end{definition}
We recall the following theorem from \cite[Theorem~3.2]{hoff11} which shows that given $\vu\in\setb$, there is a corresponding solution $\rho$ to the mass equation \eqref{Navier Stokes}$_1$ with initial data $\rho_0$.

\begin{theorem}\label{estimate on rho theorem}
Given $N>0$ and $M'_0>M_0$, there is $T_1=T_1(M_0,M'_0,N)>0$ such that if $\vu\in\setb$ for some $T>0$, then there is a unique $\rho\in\tilde\sum_1^{M'_0,\min\{T_1,T\}}$ such that the pair $(\rho,\vu)$ satisfies the equation \eqref{Navier Stokes}$_1$ such that
\begin{align}\label{bound on nabla t rho thm}
\int_0^{\min\{T_1,T\}}|\nabla\rho_t|^2dxdt\le 1
\end{align}
and 
\begin{align}\label{bound on rho t thm}
\sup_{0\le t\le\min\{T_1,T\}}\|\rho_t(\cdot,t)\|_{L^2}\le C_1M_0^2
\end{align}
for a constant $C_1>0$ which depends only on $\Omega$.  Also, there is a constant $C=C(M_0,M'_0,N)$ such that
\begin{align}\label{holder continuity of rho thm}
\langle\rho\rangle^{\frac{1}{2},\frac{1}{2}}_{\bar\Omega\times[0,T]}\le C
\end{align}
and
\begin{align}\label{continuity of rho and rhot in time thm}
\|\rho(\cdot,t_1)-\rho(\cdot,t_2)\|_{H^1},\|\rho_t(\cdot,t_1)-\rho_t(\cdot,t_2)\|_{L^2}\le C|t_2-t_1|
\end{align}
for all $t_1$, $t_2\in[0,\min\{T_1,T\}]$.
\end{theorem}

Next, we reverse the role of $\rho$ and $\vu$ and obtain estimates on $\vu$ determined by a given density $\rho\in\seta$. The results are summarised in the following lemma:

\begin{lemma}\label{estimate on u lemma}
Given $M\ge M_0$, there is a positive time $T_2 = T_2(M_0,M)$ and a constant $N = N(M_0,M)$ independent of $K$ such that if $\rho\in\seta$ and satisfies \eqref{bound on nabla t rho thm}-\eqref{continuity of rho and rhot in time thm} for some $T>0$, then there is a unique $\vu\in\tilde\sum_2^{N,\min\{T_2,T\}}$ with $\vu(\cdot,t)=\vu_0$ satisfying
\begin{align}\label{holder estimates on u in space time}
\langle\vu\rangle^{\frac{1}{2},\frac{1}{4}}_{\bar\Omega\times[0,T]}\le N,
\end{align}
\begin{align}\label{continuity in time for u and ut}
\|\vu(\cdot,t_2)-\vu(\cdot,t_1)\|_{H^2},\|\vu_t(\cdot,t_2)-\vu_t(\cdot,t_1)\|_{L^2}\le N|t_2-t_1|^\frac{1}{2},
\end{align}
for $t_1$, $t_2\in[0,\min\{T_2,T\}]$.
\end{lemma}

\begin{proof}
It suffices to obtain the bound \eqref{bound on u def} on $\vu$ as the estimates \eqref{holder estimates on u in space time}-\eqref{continuity in time for u and ut} follow by the same argument given in \cite[Lemma~3.4]{hoff11}. The proof consists of four separate energy-type estimates which will be carried out in subsequent steps. Most of the details are reminiscent of those given in \cite{hoff11} except we have to ensure that the constant $N$ is independent of $K$. Throughout this proof $N = N(M_0,M)$ will denote a generic positive constant as described in the statement of the lemma.

\noindent{\bf Step 1. Preliminary $L^2$ bound:} We multiply \eqref{Navier Stokes}$_2$ by $u^j$, apply the boundary condition \eqref{Navier condition} and sum over $j$ to obtain
\begin{align*}
&\frac{d}{dt}\intox\frac{1}{2}\rho|\vu|^2dx+\mu\intox|\nabla\vu|^2dx+K^{-1}\intoxb|\vu|^2dS_x\\
&\le N\intox(|\nabla P(\rho)||\vu|+|\rho_t+\divv(\rho\vu)||\vu|^2)dx.
\end{align*}
Since $K>0$, the boundary integral can be discarded from the left side, and hence we apply the bound \eqref{bound on rho def} on $\rho$ to conclude
\begin{align}\label{L2 bound on u}
\frac{d}{dt}\intox|\vu|^2dx+\intox|\nabla\vu|^2dx\le N(1+\|\vu(\cdot,t)\|^3_{H^1}).
\end{align}
\noindent{\bf Step 2. $H^1$ bound:} Next we multiply \eqref{Navier Stokes}$_2$ by $u^j_t$, apply the boundary condition \eqref{Navier condition} and sum over $j$ to obtain
\begin{align*}
&\frac{d}{dt}\Big[\intox\frac{1}{2}(\mu|\nabla\vu|^2+\lambda(\divv(\vu))^2dx+K^{-1}\intoxb\frac{1}{2}|\vu|^2dS_x\Big]+N^{-1}\intox|\vu_t|^2dx\\
&\le N\Big[1+\intox|\vu|^2|\nabla\vu|^2dx\Big].
\end{align*}
Using Lemma~\ref{estimate on lame operator lemma} and the embedding \eqref{embedding from Lp to H1}, we have
\begin{align}\label{L2 bound on Lu}
\intox|\cL\vu|^2dx\le N\left(\intox(|\vu_t|^2+|\rho|^2)+\intox|\vu|^2|\nabla\vu|^2\right),
\end{align}
and the last integral on the right of the above can be bounded by
\begin{align*}
\intox|\vu|^2|\nabla\vu|^2dx&\le N(\|\vu|^4_{H^1}+\|\vu\|^3_{H^1}\|\cL\vu\|_{L^2})\\
&\le N\left(\|\vu|^4_{H^1}+\|\vu\|^3_{H^1}\left(\intox(|\vu_t|^2+|\rho|^2)+\intox|\vu|^2|\nabla\vu|^2\right)^\frac{1}{2}\right),
\end{align*}
which gives
\begin{align*}
\intox|\vu|^2|\nabla\vu|^2dx\le N(1+\|\vu\|^6_{H^1}+\|\vu\|^3_{H^1}\|\vu_t\|_{L^2}).
\end{align*}
Hence there is a positive time $T_2=T_2(M_0,M)$ which is independent of $K$ such that $\vu$ satisfies
\begin{align}\label{H1 bound on u}
\sup_{0\le t\le\min\{T_2,T\}}\|\vu(\cdot,t)\|^2_{H^1}+\int_0^{\min\{T_2,T\}}\intox|\vu_t|^2dxdt\le N.
\end{align}
\noindent{\bf Step 3. $H^2$ bound:} We then multiply \eqref{Navier Stokes}$_2$ by $u^j_{tt}$, apply the boundary condition \eqref{Navier condition} and sum over $j$ to obtain
\begin{align*}
\frac{d}{dt}\intox\frac{1}{2}\rho|\vu_t|^2dx+N^{-1}\intox|\nabla\vu_t|^2dx+K^{-1}\intoxb|\vu_t|^2dS_x\le N(1+\|\vu_t\|^4_{L^2}).
\end{align*}
Again since $K>0$, the boundary integral as appeared above can be discarded. It follows that for a new time $T_2 = T_2(M_0,M)$, 
\begin{align}\label{L2 bound on ut}
\sup_{0\le t\le \min\{T_2,T\}}\intox|\vu_t|^2dx+\int_0^{\min\{T_2,T\}}\intox|\nabla\vu_t|^2dxdt\le N.
\end{align}
Furthermore, by applying the bounds \eqref{H1 bound on u} and \eqref{L2 bound on ut} on \eqref{L2 bound on Lu}, we can see that $\|\cL\vu\|_{L^2}$ can be bounded by $N$, which implies that
\begin{align}\label{H2 bound on u}
\sup_{0\le t\le\min\{T_2,T\}}\|\vu(\cdot,t)\|_{H^2}\le N.
\end{align}
\noindent{\bf Step 4. $H^3$ bound:} Finally, we multiply \eqref{Navier Stokes}$_2$ by $\cL u^j_t$ and sum over $j$ to obtain
\begin{align}\label{H3 bound step 1}
\intox\{\rho(\vu_t+\nabla\vu\,\vu)+\nabla P(\rho)-\cL\vu\}_t\cdot\cL\vu_t dx=0.
\end{align}
We compute the term $\intox\rho\vu_{tt}\cdot\Delta\vu_t$ as appeared in \eqref{H3 bound step 1}. Using the boundary condition \eqref{Navier condition},
\begin{align*}
\intox\rho\vu_{tt}\cdot\Delta\vu_t&=-\frac{d}{dt}\intox\frac{1}{2}\rho|\nabla\vu_t|^2dx+\intox\frac{1}{2}\rho_t|\nabla\vu_t|^2dx\\
&\qquad-K^{-1}\frac{d}{dt}\intoxb\rho|\vu_t|^2dS_x-\intoxb\frac{1}{2}\rho_t\vu_t\cdot\nabla\vu_tdS_x,
\end{align*}
and by applying \eqref{bound on bdry integral}, the term $-\intoxb\frac{1}{2}\rho_t\vu_t\cdot\nabla\vu_tdS_x$ can be bounded by
\begin{align*}
&\left|\intoxb\frac{1}{2}\rho_t\vu_t\cdot\nabla\vu_tdS_x\right|\\
&\le\frac{1}{2}\left(\intoxb|\rho_t\vu_t|^2\right)^\frac{1}{2}\left(\intoxb|\nabla\vu_t|^2\right)^\frac{1}{2}\\
&\le N\|\vu_t\|_{H^2}\left\{\Big(\intox|\rho_t\vu_t|^2\Big)^\frac{1}{2}+\Big(\intox|\rho_t\vu_t|^2\Big)^\frac{1}{4}\Big(\intox|\nabla\rho_t|^2|\vu_t|^2\Big)^\frac{1}{4}\right\}\\
&\qquad+N\|\vu_t\|_{H^2}\left\{\Big(\intox|\rho_t\vu_t|^2\Big)^\frac{1}{4}\Big(\intox|\rho_t|^2|\nabla\vu_t|^2\Big)^\frac{1}{4}\right\}.
\end{align*}
Hence we obtain
\begin{align}\label{H3 bound step 2}
\intox\rho\vu_{tt}\cdot\Delta\vu_t&\le-\frac{d}{dt}\left(\intox\frac{1}{2}\rho|\nabla\vu_t|^2dx+K^{-1}\intoxb\rho|\vu_t|^2dS_x\right)\notag\\
&\qquad+N\|\vu_t\|_{H^2}\left\{\Big(\intox|\rho_t\vu_t|^2\Big)^\frac{1}{2}+\Big(\intox|\rho_t\vu_t|^2\Big)^\frac{1}{4}\Big(\intox|\nabla\rho_t|^2|\vu_t|^2\Big)^\frac{1}{4}\right\}\notag\\
&\qquad+N\|\vu_t\|_{H^2}\left\{\Big(\intox|\rho_t\vu_t|^2\Big)^\frac{1}{4}\Big(\intox|\rho_t|^2|\nabla\vu_t|^2\Big)^\frac{1}{4}\right\}.
\end{align}
Notice that the since $K^{-1}\intoxb\rho|\vu_t|^2dS_x\ge0$, it can be dropped off from the analysis after we integrate over time. The other terms in \eqref{H3 bound step 1} can be treated in a similar way as we did before so that by applying \eqref{H3 bound step 2} on \eqref{H3 bound step 1}, integrating over time, using the bounds available for $\rho\in\seta$ and performing a long but straightforward sequence of estimates, we conclude that
\begin{align}\label{H3 bound step 3}
\sup_{0\le t\le\min\{T_2,T\}}\intox|\nabla\vu_t|^2dx+\int_0^{\min\{T_2,T\}}\intox|\cL\vu_t|^2dxdt\le N.
\end{align}
Together with \eqref{H2 bound on u} and the result \eqref{elliptic estimate on lame operator} obtained in Lemma~\ref{estimate on lame operator lemma}, we have
\begin{align}\label{H3 bound on u}
\sup_{0\le t\le T}(\|\vu(\cdot,t)\|^2_{H^3},\|\vu_t(\cdot,t)\|^2_{H^1})+\int_0^{\min\{T_2,T\}}\|\vu_t(\cdot,t)\|^2_{H^2}\le N,
\end{align}
which implies $\vu\in\tilde\sum_2^{N,\min\{T_2,T\}}$ as claimed.
\end{proof}

\begin{proof}[Proof of Theorem~\ref{local-in-time existence theorem}]
As mentioned before, we only consider the system (NSENC), which is the Navier-Stokes equations \eqref{Navier Stokes} with the Navier boundary condition \eqref{Navier condition}. With the help of the estimates obtained in Theorem~\ref{estimate on rho theorem} and Lemma~\ref{estimate on u lemma}, Theorem~\ref{local-in-time existence theorem} can now be proved by the method given in \cite{hoff11} and we only give a brief outline of its proof. First of all, by Theorem~\ref{estimate on rho theorem}, given a suitable velocity $\vu$, there is a corresponding density $\rho$ so that $(\rho,\vu)$ satisfies the mass equation \eqref{Navier Stokes}$_1$. Next by reversing the role of $\rho$ and $\vu$, we show that given suitably chosen $\rho$, there is a velocity $\vu$ which satisfies the momentum equation \eqref{Navier condition}$_2$. The key idea for constructing $\vu$ from $\rho$ is to apply Galerkin approximation to \eqref{Navier condition}$_2$, more precisely, given a suitable density $\rho$, if $V^n$ is the span of the first eigenfunctions of the Lam\'{e} operator $\cL$, then we can construct an approximate velocity $\vu^n:[0,T]\to V^n$ with some prepared initial data $\vu_0^n$. Lemma~\ref{estimate on u lemma} then applies to show that $\vu^n$ satisfies the bounds \eqref{bound on u def} and estimates \eqref{holder estimates on u in space time}-\eqref{continuity in time for u and ut} which are all independent of $n$. By taking $n\to\infty$, we obtain the velocity $\vu$ for the given $\rho$ and that the required bounds are retained in the limit. The argument will then be completed by combining the above constructions in an iterative process $\vu^{(k)}\mapsto\rho^{(k)}\mapsto\vu^{(k+1)}$. 
\end{proof}

\section{Convergence of smooth solutions: Proof of Theorem~\ref{main thm}}\label{convergence section}

In this section, we give the proof of Theorem~\ref{main thm} which will be carried out in a sequence of lemmas. We make use of the $K$-independent bounds obtained in Theorem~\ref{local-in-time existence theorem} in controlling both the densities and velocities. To begin with, suppose $(\trho,\tu)$ and $(\rho,\vu)$ are smooth classical solutions to the system \eqref{Navier Stokes} which are defined on $\Omega \times [0, T]$ with boundary conditions \eqref{Navier condition} and \eqref{no slip condition} respectively satisfying the bounds \eqref{boundedness on rho local-in-time solution}-\eqref{boundedness on u local-in-time solution}, and assume that $(\trho,\tu)$ and $(\rho,\vu)$ are having the same initial data $(\rho_0,\vu_0)$ which satisfy \eqref{boundedness condition on initial data}-\eqref{compatibility condition on u0 no-slip}. Define
\begin{align*}
\vw:=\vu-\tu,\qquad\phi:=\rho-\trho.
\end{align*}
Then we have $\vw\Big|_{t=0}=0$ and $\phi\Big|_{t=0}=0$. Furthermore, for all $t\in[0,T]$, $\vw$ and $\phi$ satisfy the following integral equations respectively:
\begin{align}\label{identity for phi}
\frac{1}{2}\intox|\phi|^2dx+\intoxt\phi\divv(\rho \vu-\trho\tu)dxds=0.
\end{align}
and
\begin{align}\label{energy equality for w}
&\intoxt(\rho \vu-\trho\tu)_t\cdot \vw dxds+\intoxt\divv(\rho \vu\otimes \vu-\trho\tu\otimes\tu)\cdot \vw dxds\notag\\
&\qquad+\intoxt(\nabla P(\rho)-\nabla P(\trho))\cdot \vw dxds+\mu\intox|\nabla \vw|^2+\lambda\intox(\divv (\vw))^2dxds\notag\\
&=-\intoxtb(K^{-1}\vu-\No\cdot\nabla\tu)\cdot \vw dS_xds.
\end{align}
With the help of those $K$-independent bounds obtained in Theorem~\ref{local-in-time existence theorem}, we can bound $\phi$ and $\vw$ which will be given in subsequent lemmas. 

We first prove the following lemma which gives an estimate on $\phi$:
\begin{lemma}\label{estimate on phi lemma}
For all $t\in[0,T]$, we have
\begin{align}\label{estimate on phi}
\intox\phi^2(x,t)dx&\le C\Big(\intoxtb|\vu|^2dS_xds\Big)^\frac{1}{2}+C\intoxt|\phi|^2dxds\notag\\
&\qquad+C\Big(\intoxt|\phi|^2dxds\Big)^\frac{1}{2}\Big(\intoxt|\nabla \vw|^2dxds\Big)^\frac{1}{2}\notag\\
&\qquad+C\Big(\intoxt|\phi|^2dxds\Big)^\frac{1}{2}\Big(\intoxt|\vw|^2dxds\Big)^\frac{1}{2},
\end{align}
where $C$ is a positive constant which only depends on $T$ and on $M_0$, $M_0'$ and $N$ as described in Theorem~\ref{local-in-time existence theorem}, and is independent of $K$.
\end{lemma}

\begin{proof}
In view of \eqref{identity for phi}, since $\rho \vu-\trho\tu=\phi \vu+\trho \vw$, we can decompose the integral $\dis\intoxt\phi\divv(\rho \vu-\trho\tu)dxds$ as follows:
\begin{align*}
\intoxt\phi\divv(\rho \vu-\trho\tu)dxds=\intoxt\phi\divv(\phi \vu+\trho \vw)dxds:=I_1+I_2,
\end{align*}
where 
\begin{align*}
I_1:=\intoxt\phi\divv(\phi \vu)dxds,\qquad I_2:=\intoxt\phi\divv(\trho \vw)dxds.
\end{align*}
To estimate $I_1$, we notice that
\begin{align*}
\intoxt\phi\divv(\phi \vu)=\intoxt\frac{1}{2}\divv(\phi^2 \vu)+\intoxt\frac{1}{2}\phi^2\divv(\vu),
\end{align*}
and therefore by the bounds \eqref{boundedness on u local-in-time solution} and \eqref{bound on L infty norm} and the boundary condition \eqref{Navier condition}, 
\begin{align}\label{bound on I1}
I_1&=\frac{1}{2}\intoxtb\phi^2 \No\cdot \vu dS_xds+\frac{1}{2}\intoxt\phi^2\divv(\vu)dxds\notag\\
&\le \frac{1}{2}\|\divv(\vu)\|_{L^\infty}\intoxt|\phi|^2dxds\le C\intoxt|\phi|^2dxds.
\end{align}
For the term $I_2$, using the bounds \eqref{boundedness on rho local-in-time solution}-\eqref{boundedness on u local-in-time solution} and \eqref{bound on L infty norm}, we readily have
\begin{align*}
I_2&\le C\intoxt|\trho||\nabla \vw||\phi|dxds+C\intoxt|\nabla\trho||\vw||\phi|dxds\notag\\
&\le C\|\trho\|_{L^\infty}\intoxt|\nabla \vw||\phi|dxds+C\int_0^t\|\nabla\trho\|_{L^4}\|\vw\|_{L^4}\|\phi\|_{L^2}dxds.
\end{align*}
Using the bound \eqref{boundedness on rho local-in-time solution} on $\trho$, the term $C\|\trho\|_{L^\infty}\intoxt|\nabla \vw||\phi|dxds$ can be bounded by $$C\Big(\intoxt|\phi|^2dxds\Big)^\frac{1}{2}\Big(\intoxt|\nabla \vw|^2dxds\Big)^\frac{1}{2},$$ and for $C\int_0^t\|\nabla\trho\|_{L^4}\|\vw\|_{L^4}\|\phi\|_{L^2}dxds$, we apply the embedding \eqref{embedding from Lp to H1} and the bound \eqref{boundedness on rho local-in-time solution} to obtain
\begin{align*}
&C\int_0^t\|\nabla\trho\|_{L^4}\|\vw\|_{L^4}\|\phi\|_{L^2}dxds\\
&\le C\int_0^t(\|\nabla\trho\|_{L^2}+\|\nabla\trho\|_{L^2}^\frac{1}{4}\|\nabla^2\trho\|_{L^2}^\frac{1}{4})(\|\vw\|_{L^2}+\|\vw\|_{L^2}^\frac{3}{4}\|\nabla\vw\|_{L^2}^\frac{1}{4})\|\phi\|_{L^2}\\
&\le C\Big(\intoxt|\phi|^2dxds\Big)^\frac{1}{2}\Big(\intoxt|\nabla \vw|^2dxds\Big)^\frac{1}{2}\notag\\
&\qquad+C\Big(\intoxt|\phi|^2dxds\Big)^\frac{1}{2}\Big(\intoxt|\vw|^2dxds\Big)^\frac{1}{2}.
\end{align*}
Hence we have
\begin{align}\label{bound on I2}
I_2&\le C\Big(\intoxt|\phi|^2dxds\Big)^\frac{1}{2}\Big(\intoxt|\nabla \vw|^2dxds\Big)^\frac{1}{2}\notag\\
&\qquad+C\Big(\intoxt|\phi|^2dxds\Big)^\frac{1}{2}\Big(\intoxt|\vw|^2dxds\Big)^\frac{1}{2}.
\end{align}
We apply the bounds \eqref{bound on I1} and \eqref{bound on I2} on \eqref{identity for phi}, and the assertion \eqref{estimate on phi} follows.
\end{proof}
Next we prove the following lemma which consists of the estimate on $\vw$:
\begin{lemma}\label{estimate on w lemma}
For all $t\in[0,T]$, we have
\begin{align}\label{estimate on w}
&\intox\rho|\vw(x,t)|^2dx+\intoxt|\nabla \vw|^2dxds\notag\\
&\le C\intoxt|\vw|^2dxds+C\Big(\intoxt|\vw|^2dxds\Big)^\frac{1}{2}\Big(\intoxt|\nabla \vw|^2dxds\Big)^\frac{1}{2}\notag\\
&\,\,\,+C\Big(\intoxt|\nabla \vw|^2dxds\Big)^\frac{1}{2}\Big(\intoxt|\phi|^2dxds\Big)^\frac{1}{2}\notag\\
&\,\,\,+C(\|\tu_t\|_{H^2}+1)\Big(\intoxt|\vw|^2dxds\Big)^\frac{1}{2}\Big(\intoxt|\phi|^2dxds\Big)^\frac{1}{2}\notag\\
&\,\,\,+C\intoxtb|\vu|^2dS_xds+C\Big(\intoxtb|\vu|^2dS_xds\Big)^\frac{1}{2},
\end{align}
where $C$ is a positive constant which only depends on $T$ and on $M_0$, $M_0'$ and $N$ as described in Theorem~\ref{local-in-time existence theorem}, and is independent of $K$.
\end{lemma}

\begin{proof}
In view of \eqref{energy equality for w}, we first estimate the left side of \eqref{energy equality for w}. Define
\begin{align*}
I_3&:=\intoxt(\rho \vu-\trho\tu)_t\cdot w dxds,\qquad I_4:=\intoxt\divv(\rho \vu\otimes \vu-\trho\tu\otimes\tu)\cdot \vw dxds\\
I_5&:=\intoxt(\nabla P(\rho)-\nabla P(\trho))\cdot \vw dxds,\qquad I_6:=\mu\intox|\nabla \vw|^2+\lambda\intox(\divv (\vw))^2dxds.
\end{align*}
To estimate $I_3$, we note that
\begin{align*}
&\intoxt(\rho \vu-\trho\tu)_t\cdot \vw dxds\\
&=\intoxt\frac{1}{2}(\rho|\vw|^2)_tdxds+\intoxt\frac{1}{2}\rho_t|\vw|^2dxds+\intoxt(\phi\tu)_t\cdot \vw dxds,
\end{align*}
hence
\begin{align}\label{eqn for I3}
I_3=\frac{1}{2}\intox\rho|\vw|^2dx+I_{3,1}+I_{3,2}+I_{3,3}
\end{align}
where
\begin{align*}
I_{3,1}:=\intoxt\frac{1}{2}\rho_t|\vw|^2dxds,\,\,\, I_{3,2}:=\intoxt\phi_t\tu\cdot \vw dxds,\,\,\,I_{3,3}:=\intoxt\phi\tu_t\cdot \vw dxds.
\end{align*}
The term $I_{3,3}$ is readily bounded by
\begin{align*}
|I_{3,3}|\le C\|\tu_t\|_{H^2}\Big(\intoxt|\vw|^2dxds\Big)^\frac{1}{2}\Big(\intoxt|\phi|^2dxds\Big)^\frac{1}{2}.
\end{align*}
For $I_{3,1}$, we use the mass equation \eqref{Navier Stokes}$_1$ for $(\rho,\vu)$ and notice that $\vw=\vu$ and $\No\cdot \vu=0$ on $\partial\Omega$, we have, by the bounds \eqref{boundedness on rho local-in-time solution}-\eqref{boundedness on u local-in-time solution} and the estimate \eqref{bound on L infty norm} on $\|\rho \vu\|_{L^\infty}$ that
\begin{align*}
|I_{3,1}|&\le C\intoxt|\rho \vu||\nabla \vw||\vw|dxds\\
&\le C\|\rho \vu\|_{L^\infty}\Big(\intoxt|\vw|^2dxds\Big)^\frac{1}{2}\Big(\intoxt|\nabla \vw|^2dxds\Big)^\frac{1}{2}\\
&\le C\Big(\intoxt|\vw|^2dxds\Big)^\frac{1}{2}\Big(\intoxt|\nabla \vw|^2dxds\Big)^\frac{1}{2},
\end{align*}
while for the term $I_{3,2}$, we notice that 
\begin{align*}
&\intoxt\phi_t\tu\cdot \vw dxds\\
&=-\intoxt\divv(\rho \vu-\trho\tu)\tu\cdot \vw dxds=-\intoxt\divv(\phi \vu+\trho \vw)\tu\cdot \vw dxds
\end{align*}
and recall the fact that $\tu=0$ on $\partial\Omega$ to obtain
\begin{align*}
|I_{3,2}|&=\Big|\intoxt\divv(\rho \vu-\trho\tu)\tu\cdot \vw dxds\Big|= \Big|\intoxt(\phi \vu+\trho \vw)\cdot\nabla(\tu\cdot \vw)dxds\Big|,
\end{align*}
which gives
\begin{align*}
|I_{3,2}|&\le C\intoxt|\vw|^2dxds+C\Big(\intoxt|\vw|^2dxds\Big)^\frac{1}{2}\Big(\intoxt|\nabla \vw|^2dxds\Big)^\frac{1}{2}\\
&\qquad+C\Big(\intoxt|\vw|^2dxds\Big)^\frac{1}{2}\Big(\intoxt|\phi|^2dxds\Big)^\frac{1}{2}\\
&\qquad+C\Big(\intoxt|\nabla \vw|^2dxds\Big)^\frac{1}{2}\Big(\intoxt|\phi|^2dxds\Big)^\frac{1}{2}.
\end{align*}
Hence we can bound $I_{3,1}+I_{3,2}+I_{3,3}$ by
\begin{align}\label{bound on I3s}
&|I_{3,1}+I_{3,2}+I_{3,3}|\notag\\
&\le C\intoxt|\vw|^2dxds+C\Big(\intoxt|\vw|^2dxds\Big)^\frac{1}{2}\Big(\intoxt|\nabla \vw|^2dxds\Big)^\frac{1}{2}\notag\\
&\,\,\,+C\Big(\intoxt|\nabla \vw|^2dxds\Big)^\frac{1}{2}\Big(\intoxt|\phi|^2dxds\Big)^\frac{1}{2}\notag\\
&\,\,\,+C(\|\tu_t\|_{H^2}+1)\Big(\intoxt|\vw|^2dxds\Big)^\frac{1}{2}\Big(\intoxt|\phi|^2dxds\Big)^\frac{1}{2}.
\end{align}
Next we consider $I_4$. Upon integrating by parts, using the boundary condition that $\No\cdot \vu=0$ on $\partial \Omega$ and applying bounds \eqref{boundedness on rho local-in-time solution}-\eqref{boundedness on u local-in-time solution}, it can be estimated as follows (summation over repeated indexes is understood):
\begin{align}\label{bound on I4}
|I_4|&\le \Big|\intoxt\nabla w^j\cdot(\rho u^j \vu-\trho\tu^j \tu)dxds\Big|\notag\\
&\le C\Big(\intoxt|\nabla \vw|^2dxds\Big)^\frac{1}{2}\Big(\intoxt|\phi|^2dxds\Big)^\frac{1}{2}\notag\\
&\qquad+C\Big(\intoxt|\nabla \vw|^2dxds\Big)^\frac{1}{2}\Big(\intoxt|\vw|^2dxds\Big)^\frac{1}{2}.
\end{align}
For the term $I_5$, we again integrate by parts and use the boundary condition to obtain
\begin{align}\label{bound on I5}
|I_5|&\le\Big|\intoxtb \No\cdot \vu(P(\rho)-P(\trho))dS_xds\Big|+\Big|\intoxt(P(\rho)-P(\trho))\cdot\nabla \vw dxds\Big|\notag\\
&\le C\Big(\intoxt|\nabla \vw|^2dxds\Big)^\frac{1}{2}\Big(\intoxt|\rho-\trho|^2dxds\Big)^\frac{1}{2}.
\end{align}
Now we estimate the right side of \eqref{energy equality for w}. Recalling that $\vw=\vu$ on $\partial\Omega$ and applying the estimate \eqref{bound on bdry integral} on $\nabla \tu$ to get
\begin{align}\label{bound on bdry terms}
&-\mu\intoxtb(K^{-1}\vu-\No\cdot\nabla\tu)\cdot \vw dS_xds\notag\\
&=-\mu\intoxtb K^{-1}|\vu|^2dS_xds+\mu\intoxtb (\No\cdot\nabla\tu)\cdot \vu dS_xds\notag\\
&\le C\Big(\intoxtb|\nabla\tu|^2dS_xds\Big)^\frac{1}{2}\Big(\intoxtb|\vu|^2dS_xds\Big)^\frac{1}{2}\notag\\
&\le C\Big(\intoxtb\{|\nabla\tu|^2+|\nabla\tu||\Delta\tu|\}dS_xds\Big)^\frac{1}{2}\Big(\intoxtb|\vu|^2dS_xds\Big)^\frac{1}{2}\notag\\
&\le C\Big(\intoxtb|\vu|^2dS_xds\Big)^\frac{1}{2}.
\end{align}
Combining \eqref{eqn for I3}, \eqref{bound on I3s}, \eqref{bound on I4}, \eqref{bound on I5}, \eqref{bound on bdry terms} with \eqref{energy equality for w}, we conclude that
\begin{align*}
&\intox\rho|\vw|^2dx+\intoxt|\nabla \vw|^2dxds\\
&\le C\intoxt|\vw|^2dxds+C\Big(\intoxt|\vw|^2dxds\Big)^\frac{1}{2}\Big(\intoxt|\nabla \vw|^2dxds\Big)^\frac{1}{2}\notag\\
&\,\,\,+C(\|\tu_t\|_{H^2}+1)\Big(\intoxt|\vw|^2dxds\Big)^\frac{1}{2}\Big(\intoxt|\phi|^2dxds\Big)^\frac{1}{2}\notag\\
&\,\,\,+C\Big(\intoxt|\nabla \vw|^2dxds\Big)^\frac{1}{2}\Big(\intoxt|\phi|^2dxds\Big)^\frac{1}{2}+C\Big(\intoxtb|\vu|^2dS_xds\Big)^\frac{1}{2}
\end{align*}
and the estimate \eqref{estimate on w lemma} follows.
\end{proof}
The following lemma contains the crucial bound on $\|\vu\|_{L^2(\partial\Omega)}$ in terms of $K$ which will be used for proving Theorem~\ref{main thm}.
\begin{lemma}\label{estimate on u lemma}
There exists $T^*\in(0,T]$ such that for all $t\in[0,T^*]$, we have
\begin{align}\label{L2 bound on u on bdry}
\intoxtb |\vu|^2 dS_xds\le M_0K.
\end{align}
\end{lemma}

\begin{proof}
We multiply the momentum equation \eqref{Navier Stokes}$_2$ by $u^j$, sum over $j$ and integrate to obtain
\begin{align}\label{equality from momentum eqn}
&\intox\frac{\rho|\vu|^2}{2}dx+\intoxt \vu\cdot\nabla Pdxds+\intoxt\{\mu|\nabla \vu|^2+\lambda(\divv(\vu))^2\}dxds\notag\\
&=\intox\frac{\rho_0|\vu_0|^2}{2}dx+\intoxtb\mu(\nabla \vu\No)\cdot \vu dS_xds+\intoxtb\lambda\divv(\vu)\vu\cdot \No dS_xds\notag\\
&=\intox\frac{\rho_0|\vu_0|^2}{2}dx-\intoxtb K^{-1}|\vu|^2 dS_xds,
\end{align}
where the last equality of \eqref{equality from momentum eqn} follows from the boundary condition \eqref{Navier condition}. Upon integrating by parts and using the boundary condition that $\No\cdot \vu=0$ on $\partial\Omega$, we have
\begin{align*}
\Big|\intoxt \vu\cdot\nabla Pdxds\Big|&\le \Big(\intoxt|\nabla \vu|^2dxds\Big)^\frac{1}{2}\Big(\intoxt|P|^2dxds\Big)^\frac{1}{2}\\ 
&\le C\Big(\intoxt|\nabla \vu|^2dxds\Big)^\frac{1}{2}\Big(\intoxt|\rho|^2dxds\Big)^\frac{1}{2},
\end{align*}
and hence we obtain from \eqref{equality from momentum eqn} that
\begin{align}\label{L2 estimate on u}
&\intox\frac{\rho|\vu|^2}{2}dx+\intoxt\{\mu|\nabla \vu|^2+\lambda(\divv(\vu))^2\}dxds\notag\\
&\le C\Big(\intoxt|\nabla \vu|^2dxds\Big)^\frac{1}{2}\Big(\intoxt|\rho|^2dxds\Big)^\frac{1}{2}+\intox\frac{\rho_0|\vu_0|^2}{2}dx\notag\\
&\qquad-\intoxtb K^{-1}|\vu|^2 dS_xds.
\end{align}
On the other hand, we make use of the boundary condition \eqref{Navier condition} and the bound \eqref{boundedness on rho local-in-time solution} on $\rho$ to obtain
\begin{align}\label{L2 estimate on rho}
\intox\frac{|\rho|^2}{2}dx&=\intox\frac{|\rho_0|^2}{2}dx-\intoxt\rho\divv(\rho \vu)dxds\notag\\
&\le \intox|\rho_0|^2dx+\intoxt|\nabla\rho||\rho||\vu|dxds\notag\\
&\le \intox|\rho_0|^2dx+C\Big(\intoxt|\rho|^2dxds\Big)^\frac{1}{2}\Big(\intoxt|\vu|^2dxds\Big)^\frac{1}{2}.
\end{align}
We sum up \eqref{L2 estimate on u} and \eqref{L2 estimate on rho}, recall the positive lower bound \eqref{boundedness on rho local-in-time solution} on $\rho$ and choose $T^*\in(0,T]$ small enough, then for $t\in[0,T^*]$, we have
\begin{align}\label{L2 estimate on rho and u}
&\sup_{s\in[0,t]}\Big(\intox|\rho|^2dx+\intox|\vu|^2dx\Big)\notag\\
&\le\intox\frac{\rho_0|\vu_0|^2}{2}dx+\intox\frac{|\rho_0|^2}{2}dx-\intoxtb K^{-1}|\vu|^2 dS_xds.
\end{align}
We deduce from \eqref{L2 estimate on rho and u} that
\begin{align*}
\intoxtb |\vu|^2 dS_xds\le M_0K,\qquad\mbox{for all $t\in[0,T^*]$,}
\end{align*}
which implies \eqref{L2 bound on u on bdry}.
\end{proof}
We are now ready to give the proof of Theorem~\ref{main thm}:
\begin{proof}[Proof of Theorem~\ref{main thm}]
We sum up \eqref{estimate on phi} and \eqref{estimate on w}, apply the positive lower bound \eqref{boundedness on rho local-in-time solution} on $\rho$ and apply Young's inequality to obtain, for $t\in[0,T]$,
\begin{align}\label{L2 bound on phi and w}
&\intox\{\phi^2(x,t)+|\vw|^2(x,t)\}dx+\intoxt|\nabla \vw|^2dxds\notag\\
&\le C(\|\tu_t\|_{H^2}+1)\intoxt\{\phi^2+|\vw|^2\}dxds\notag\\
&\qquad+C\intoxtb|\vu|^2dS_xds+C\Big(\intoxtb|\vu|^2dS_xds\Big)^\frac{1}{2}.
\end{align}
Applying Gr\"{o}nwall's inequality on \eqref{L2 bound on phi and w} and using the bound \eqref{boundedness on u local-in-time solution} on the time integral $\int_0^T\|\tu_t(\cdot,t)\|_{H^2}^2dt$, we further get
\begin{align}\label{estimate on L2 norm of phi and w step 1}
&\intox\{\phi^2(x,t)+|\vw|^2(x,t)\}dx\notag\\
&\le C\Big\{\intoxtb|\vu|^2dS_xds+\Big(\intoxtb|\vu|^2dS_xds\Big)^\frac{1}{2}\Big\}e^{Ct},\,\,\,\mbox{for all $t\in[0,T]$.}
\end{align}
Let $T^*>0$ be chosen as in Lemma~\ref{estimate on u lemma}. We apply \eqref{L2 bound on u on bdry} on \eqref{estimate on L2 norm of phi and w step 1} to obtain, for all $t\in[0,T^*]$, 
\begin{align}\label{estimate on L2 norm of phi and w step 2}
\intox\{\phi^2(x,t)+|\vw|^2(x,t)\}dx\le C\Big\{M_0K+M_0^\frac{1}{2}K^\frac{1}{2}\Big\}e^{Ct}.
\end{align}
Hence by taking $K\to0$ in \eqref{estimate on L2 norm of phi and w step 2}, we have that
\begin{align}\label{convergence of phi and u in L2}
\left\{ \begin{array}{l}
\rho\to\trho\mbox{ in $L^\infty([0,T^*];L^2(\Omega))$ as $K\to0$,} \\
\vu\to\tu\mbox{ in $L^\infty([0,T^*];L^2(\Omega))$ as $K\to0$.}
\end{array}\right.
\end{align}
Moreover, by applying the convergences given in \eqref{convergence of phi and u in L2} on \eqref{estimate on L2 norm of phi and w step 2}, we have 
\begin{align*}
\vu\to\tu\mbox{ in $L^2([0,T^*];H^1(\Omega))$ as $K\to0$.}
\end{align*}
Since $\tu=0$ on $\partial\Omega$, using \eqref{L2 bound on u on bdry} from Lemma~\ref{estimate on u lemma}, we conclude that $\vu\to\tu$ in $L^2([0,T^*];L^2(\partial\Omega))$ as $K\to0$ and \eqref{convergence of u on the boundary} follows. Finally, by Sobolev inequality, for $s_1\in(0,2]$, there exists $\sigma=\sigma(s_1)\in(0,1)$ such that for $t\in[0,T^*]$,
\begin{align}\label{Hs estimate on u}
\|(\vu-\tu)(\cdot,t)\|_{H^{s_1}}\le\|(\vu-\tu)(\cdot,t)\|_{L^2}^\sigma\|(\vu-\tu)(\cdot,t)\|_{H^{3}}^{1-\sigma},
\end{align}
so that by applying the bound \eqref{boundedness on u local-in-time solution} on $\|(\vu-\tu)(\cdot,t)\|_{H^{3}}^{1-\sigma}$, we have 
\begin{align*}
\mbox{$\dis\sup_{0\le t\le T^*}\|(\vu-\tu)(\cdot,t)\|_{H^{s_1}}\to0$ as $K\to0$}
\end{align*}
and \eqref{convergence of u in Hs} follows. The proof of \eqref{convergence of rho in Hs} is just similar and we finish the proof of Theorem~\ref{main thm}.
\end{proof}



\end{document}